UT Dec 23 2021 - Split Night INAF / LBTO-DEC



Night Info

  • Sunset 17:23
  • Twlight Ends 18:46
  • Moonrise 19:27
  • Night Center 00:18
    • If handing off between bent Gregorian foci - Handover 00:12 (12min split between partners)
    • if handing off to LBCs (or LBCs to PEPSI-Pol) - Handover 00:07 (22 min split between partners)
  • Twlight Begins 05:50

Night Plan

  • If PEPSI-Pols have been installed and connected
    • Any seeing - up to 1-2mags cloud extinction
    • AGW4 probe alignment to rotator center (1 hr) - J.Hill
    • Active Zernike angle adjustments IT 7545 (1-2 hr) - J.Hill
    • PEPSIPOL ADC debug/calibration (1 hr) - I.Ilyin
    • If time is available after PEPSI-Pol work done
      • < 2.5 hours left
        • GRB 21121A LUCI-2 imaging if conditions are good - 1.25hr w/overheads
          • /home/lbto/Service_Observing/2021B_D_E/LUCI/GRB211211A/Grb211211A_LD2021B-002_OT.xml
          • Seeng <1.2", thin clouds
          • Visible from 04hr onwards (end of night)
          • Ks-band 1hr - or - J-band 1hr - coordinate with MMT regarding which filter to attempt
        • Local Merger Backup programs if conditions are poor
      • if >2.5 hours left
        • TMS/LBC observations - any conditions
          • Lower priority than GRB in good conditions - TMS/LBC data needed more for start of night rather than end of night
  • If PEPSI-Pols have NOT been installed and connected
    • LBC TMS Testing - half night
      • TMS Fields:
  • If PEPSI-Pols have NOT been install and connected - AND LBC TMS cannot be executed
    • If conditions are good (seeing < 1", thin clouds)
      • GRB 211211A - LUCI-2 imaging
        • /home/lbto/Service_Observing/2021B_D_E/LUCI/GRB211211A
        • Ks-band 1hr - or - J-band 1hr
        • Visible from 04hr onwards (end of night)
  • Poor Condition Backup - to be used as needed during night
    • Local Mergers - LUCI-2 - Rothberg
      • OT: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V34.xml
      • finder charts are in LUCI/ sub-directory
      • 300sec x 5 dithers - for monocular - repeat script at least twice - longer if clouds really thick or seeing really bad (or both)
        • Arp 156 - 23hr-twilight
          • different telluric depending on when object is observed - noted in OT
        • NGC 4004 - 02:30hr-twilight
        • NGC 3597 -04hr-twlight - low elevation (-23 DEC)

Summary

There were clouds, thick most of the time, but bright stars could be seen in the polarimeter until the very end of the night when the clouds became too thick. Soon after we opened, the clouds were thin enough that we could guide and WFS. John checked the guiding hotspot positions for both sides - these were OK. John also checked the pupil positions - for SX unchanged but for DX where it changed, he did some work to to adjust it. He wanted to work on getting active optics to converge with Z9 and Z10 on, but could not complete this work because the clouds became too thick and active presets would fail. Ilya worked on aligning the PEPSI ADC. He sent active or acquire presets to bright stars at low elevation, positioned them near but not on the pinhole where he could see the light reflected from the mirror around the fiber and then adjusted the ADC orientation and separation. He got some work done, but had not completed it when the clouds became too thick around 12-deg twilight. John did not do the alignment of the AGW4 probe to the rotator center.

It is safe to say that while some work was done, more time will be needed on Sunday to complete the PEPSI-POL checkout.

Night Log

07:20 Reconfigured to PEPSI-POL

09:46 There has been a hole for a while, so we are opening.

Initial checkout on-sky with PEPSIPOL

09:50 Checking & correcting pointing

10:02 Collimating - John set PSF "Active Optics" gains for Z9 and Z10 on both sides to 0.0, so we don't have to worry about runaways while doing other work.

10:03 Moving to a fainter star (13.8 mag) to see the sky background in the lenslets. John set active optics gains for Z9 and Z10 to 0 to prevent runaway. John said that guiding the hotspot looks OK and he made a first adjustment to the right pupil. The left pupil position should not have changed, but the right pupil was expected to have changed. Mike did some alignment of AGW4 in the lab, and John is working on fine-tuning it using GCS & IDL.

Conclusion 1: All the PEPSIPOL and AGW3/4 machinery (hardware and software) seems to be working OK. (except Z9/Z10 active optics)

10:20 John installed the first estimate of AGW4 on-axis pupil position in PEPSIPOL_R.cfg: 163.6,217.4. This measurement from one wfsc image is close to the final value, but see the refinement below.

10:22 Next up: ADC alignment. In parallel with the ADC work, John will try to get active optics to converge with Z9 and Z10 (which didn't work because of clouds).

10:25 But - it got cloudy enough that the background is more easily visible in the lenslets - John is continuing to work on the hotspot adjustments.

Conclusion 2: The guiding hotspots for AGW3 and AGW4 are correct (no changes were made). I saw a peculiarity where the AGW4 guiding hotspot looks perfect for the first couple frames after the preset, but then seems to drift off toward the bottom of the lenslet aperture by ~3 pixels. This was prior to Ilya starting to use the hotspot to adjust the star on the fiber aperture. I'm not ready to make an IT yet, but this is a strange thing that needs to be investigated. Could there be a loose mirror in the AGW with some hysteresis? Is there something funny about guiding? I noticed this twice between approx 10:30 and 11:00 UT.

AGW On-axis Pupil Position

In parallel to Ilya's work at low elevation, John was analyzing the on-axis wfs pupils we took at higher elevation before 10:25 UT. The images analyzed were 000090 - 000115 on both sides. Analysis files (for both sides) are in /home/lbto/data/20211223_PPOL/ . The "new" thing I learned (or had previously forgotten) was that the pupil position measured by IDL (wfsc_process_file, /SHOW_PUPIL) is slightly dependent on the initial hotspot value in agw_define.pro. Thus, it required one pass of iteration to get a good measurement. The First Pass of the measurement had Old Position in agw_define.pro, and the Second Pass of the measurement had First Pass in agw_define.pro, etc. These wfsc images were post-processed with scripts that called:
/lbt/lbto/supportscripts/ObserverSupport/agw_wfsc_process.py -n -f LPOL -p /Repository/20211223/GCS/ -t left_wfscimage000090.fits -l 20211223.sx.pupil.log
for each image. The measured pupil positions in the resulting log files were averaged with awk:
awk '{s+=$5; ss+=$5^2} END{print m=s/NR, sqrt(ss/NR-m^2)}' 20211223.dxc.pupil.log

DX Pupil Position hotspot_x hotspot_y rms x rms y
Old Position 145.55 206.63    
First Pass 159.76 217.45 3.03 0.55
Second Pass 161.06 217.80 2.81 0.54
Third Pass 161.06 217.80 2.81 0.54

SX Pupil Position hotspot_x hotspot_y rms x rms y
Old Position 161.55 230.41    
First Pass 163.00 231.80 0.52 0.28
Second Pass 163.00 231.80 0.52 0.28

Conclusion 3: On-axis Pupil Positions (WFS_hotspot_x, WFS_hotspot_y) have been correctly set.

The wfs hotspots (on-axis pupil positions) were updated in both IDL agw_define.pro and GCS PEPSIPOL_?.cfg.

PEPSIPOL ADC debug / calibration

10:27 Now Ilya is sending a preset to a low elevation (42 deg) star (6.7 mag; HIP 45975) for the ADC alignment. The guide star is not found on the SX side. It was found on the DX side, but Ilya does not see the on-axis target. Guide star - target separaton 4.2 arcmin. Finally see the on-axis star on DX, but still there is no guide star and no on-axis star on SX.

10:37 David is checking pointing on a brighter star. Now we see the guide star and the target star on both sides.

10:49 Ilya is going to a lower elevation star, beta Crt (V=4.46), at 30 deg elevation. Neither guide star came up (13th mag) due to clouds, but the target is seen on the DX side and is just to the upper left.

Finally, with a successful active preset to a pointing star (WT10_33 = HD95680), Ilya worked on ADC alignment.

11:19 Now slewing to 12 Cam (6.13 mag). Clouds are getting thicker.

12:32 This star is reaching 26 deg - going to another.

13:34 David closed up due to impending dawn.

-- BarryRothberg - 20 Dec 2021
Topic revision: r11 - 23 Dec 2021, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback